OpenStudy (anonymous):

A bicycle tire has a diameter of twenty-six inches. If a pebble is stuck in the tire's treads, about how far will the pebble travel with each rotation of the tire? forty point eight two inches eighty-one point six four inches five hundred thirty point six six inches two thousand one hundred twenty-two point six four inches

OpenStudy (lucaz):

the question is about the circunference of the tire C=2*pi*radius, radius = diameter/2, radius=26/2=13 C=2*pi*13 C=26pi assuming pi is 3.14 C=26*3.14=81.64 in